31. Follow-up What is the final momentum of each canoe if they are pushed apart with a force of 56 N for 1.3 s?
Answer #1

According to newton's second law

Force = impulse

F = dp/dt

dp = F*dt

= 56*1.3 = 72.8 kg m/s

so the final momentum of each canow is 72.8 kg m/s